# Forecasting realized volatility in the stock market: a path-dependent perspective

### v1

@@ Abstract
Volatility forecasting in financial markets is a topic that has received more attention from scholars. In this paper, we propose a new volatility forecasting model that combines the heterogeneous autoregressive (HAR) model with a family of path-dependent volatility models (HAR-PD). The model utilizes the long- and short-term memory properties of price data to capture volatility features and trend features. By integrating the features of path-dependent volatility into the HAR model family framework, we develop a new set of volatility forecasting models. And, we propose a HAR-REQ model based on the empirical quartile as a threshold, which exhibits stronger forecasting ability compared to the HAR-REX model. Subsequently, the predictive performance of the HAR-PD model family is evaluated by statistical tests using data from the Chinese stock market and compared with the basic HAR model family. The empirical results show that the HAR-PD model family has higher forecasting accuracy compared to the underlying HAR model family. In addition, robustness tests confirm the significant predictive power of the HAR-PD model family.
@@ 1 Introduction
By leveraging the entire historical trajectory of the underlying asset prices, the PDV model can more comprehensively capture long-term memory effects and complex dynamics, and remain highly sensitive to new market environments. The PDV model maintains forecast accuracy at different time scales and has a high sensitivity to extreme market events, which is particularly important for predicting potential large fluctuations. Parent (2022) integrated the concept of path dependence with the rough Heston model (Euch and Rosenbaum, 2018) to introduce the rough path-dependent volatility model. This model can capture a stronger Zumbach effect  [^1] The strong Zumbach effect refers to the phenomenon where the distribution of future realized volatility depends not only on past squared returns but also on the entire trajectory of past realized volatility. The weak Zumbach effect refers to the covariance between past squared returns and future realized volatility (within a given duration) being greater than the covariance between past realized volatility and future squared returns.. These research findings not only enrich the theoretical foundation of PDV models but also open up new avenues for their practical application in finance.
@@
    Aiming to enhance predictive accuracy, we have refined the HAR-REX model by substituting the conventional cumulative normal distribution threshold with an innovative approach that employs empirical quartiles to re-segment realized volatility. Our empirical analyses reveal that the revamped HAR-REQ model, along with its extension, the HAR-PD-REQ model, exhibit significantly improved forecasting efficacy across daily, weekly, and monthly volatility horizons.
@@
The remainder of this paper is structured as follows. In Section 2, we introduce the fundamental theories of volatility and the construction of volatility forecasting models. In Section 3, we present all the basic models used in this study and the methods for combining models. In Section 4, we discuss the volatility of the Chinese financial market from a path-dependent perspective and conduct an empirical analysis of the dataset. We use various statistical testing methods, including the Model Confidence Set (MCS) test (Hansen et al., 2011) and the out-of-sample $R^{2}$ test (Campbell and Thompson, 2008), to evaluate the predictive capabilities of the HAR model family and the HAR-PD model family, and we report and discuss the results. In Section 5, we present the conclusion that the proposed HAR-PD model family demonstrates superior predictive ability.
@@ 2 Theoretical framework
[…]
@@ 2.1 The PDV model
The PDV model assumes that the asset price $S_{t}$ follows a functional SDE [^2] The well-posedness of (3) has been extensively studied and is beyond the scope of this paper; interested readers may refer to Ning and Wu (2021) and Cozma and Reisinger (2018):
@@
where $\mu_{t}$ denotes a progressively measurable stochastic process, $W_{t}$ is a standard Brownian motion, and $S_{u}(u\leq t)$ is the trajectory of $S_{u}$ for 00 to $t$, representing a functional mapping from the sample space $\mathbb{R}$ to the path space $\mathcal{C}(\mathbb{R})$. In contrast to classical jump-diffusion model (1), the PDV models (3) postulate that the volatility is dependent upon the entire trajectory of asset $S_{u}(u\leq t)$ instead of $S_{t}$. This implies that the volatility is determined not only by the current price level but also by the path through which the price reaches its current level. The path-dependent variable $G_{t}$ can incorporate various metrics, including the moving average ($S^{\Delta}_{t}$), exponentially weighted moments ($Y^{\Delta}_{t}$), and historical maximum or minimum prices ($Max^{\Delta}_{t}$ and $Min^{\Delta}_{t}$), where:
[…]
@@
where $K_{\lambda_{1}}$ is a kernel function that assigns different weights to past daily returns based on the lag term $t-i$. Moreover, $K_{\lambda_{1}}$ is specified as exponential kernels:
@@ 2.2.1 Continuous and discrete components
where $\mu_{1}=\sqrt{2/\pi}$, and the factor $n/(n-1)$ is used to adjust for sample size. Barndorff-Nielsen and Shephard (2003) demonstrated that under the assumption that the price process $S_{t}$ is a semimartingale, $\mathrm{RBV}_{t}$ serves as a consistent estimator of the integrated volatility and the difference between ${\rm RV}_{t}$ and $\mathrm{RBV}_{t}$ provides an estimator for jump variation.
[…]
@@ 3.1 Benchmark: HAR model family
The model comprises three explanatory variables: $\mathrm{{RV}}_{t-1}$, $\mathrm{\overline{RV}}_{t-5}$, and $\mathrm{\overline{RV}}_{t-22}$, which correspond to short-term (daily), medium-term (weekly), and long-term (monthly) volatility, respectively. The coefficients $\beta_{0},\beta_{1}$, $\beta_{2}$, $\beta_{3}$ in (7) can be estimated using the loss function MSE for estimation. The lagged terms $\mathrm{\overline{RV}}_{t-5}$ and $\mathrm{\overline{RV}}_{t-22}$ represent 5-step and 22-step lags, respectively, and can be generalized as $\mathrm{\overline{RV}}_{t-h}$, where $h$ denotes the number of lag terms, calculated as:
@@
$\begin{split}\mathrm{RV}_{t}=\beta_{0}+\beta_{1}\mathrm{REX}_{t-1}^{+}+\beta_{2}\mathrm{\overline{REX}}_{t-5}^{+}+\beta_{3}\mathrm{\overline{REX}}_{t-22}^{+}+\beta_{4}\mathrm{REX}_{t-1}^{-}+\beta_{5}\mathrm{\overline{REX}}_{t-5}^{-}+\\ \quad\beta_{6}\mathrm{\overline{REX}}_{t-22}^{-}+\beta_{7}\mathrm{REX}_{t-1}^{m}+\beta_{8}\mathrm{\overline{REX}}_{t-5}^{m}+\beta_{9}\mathrm{\overline{REX}}_{t-22}^{m}+\epsilon_{t}\,,\end{split}$ (9)
@@
$\begin{split}\mathrm{RV}_{t}=&\beta_{0}+\beta_{1}\mathrm{REQ}_{t-1}^{+}+\beta_{2}\mathrm{\overline{REQ}}_{t-5}^{+}+\beta_{3}\mathrm{\overline{REQ}}_{t-22}^{+}+\beta_{4}\mathrm{REQ}_{t-1}^{-}+\beta_{5}\mathrm{\overline{REQ}}_{t-5}^{-}+\\ &\quad\beta_{6}\mathrm{\overline{REQ}}_{t-22}^{-}+\beta_{7}\mathrm{REQ}_{t-1}^{m}+\beta_{8}\mathrm{\overline{REQ}}_{t-5}^{m}+\beta_{9}\mathrm{\overline{REQ}}_{t-22}^{m}+\epsilon_{t}\,,\end{split}$ (10)
@@ 3.2 Encoding path-dependent: HAR-PD models
$\mathrm{RV}_{t}=\beta_{0}+\beta_{1}{R}_{2,t-1}+\beta_{2}{\overline{R}}_{2,t-5}+\beta_{3}{\overline{R}}_{2,t-22}+\epsilon_{t}\,,$
where ${R}_{2,t-1}$, ${\overline{R}}_{2,t-5}$, and ${\overline{R}}_{2,t-22}$ represent volatility features based on the moments ${t-1}$, ${t-5}$, and ${t-22}$, respectively. The construction method for ${\overline{R}}_{2,t-h}$ is similar (5).
@@
In the HAR-PD-CJ model, the volatility characteristics incorporate both jump and continuous components. The volatility characteristics describe the overall fluctuation within a specific time period, while jumps in financial markets indicate significant changes in asset prices. By incorporating jump characteristics, volatility predictions can be improved, particularly for fluctuations caused by public events, announcements, and other market uncertainties. Here, $\mathrm{{PDCJ}}_{t-1}$, $\mathrm{\overline{{PDCJ}}}_{t-5}$, and $\mathrm{\overline{{PDCJ}}}_{t-22}$ represent the jump components based on the path dependence at time $t-1$, $t-5$, $t-22$ respectively. Combined with the constructor of (4), the $\mathrm{PDCJ}_{t}$ can be constructed as:
@@
The basic HAR-RS model decomposes the realized semivariance into positive and negative components for volatility prediction. This decomposition captures both the actual price skewness within specific time periods and the dynamic evolution of realized skewness, thereby providing insights into volatility patterns. The construction method for $\mathrm{PDRS}_{t}$ is shown in (12), and the construction methods for $\mathrm{\overline{{PDRS}}}_{t-h}^{+}$ and $\mathrm{\overline{{PDRS}}}_{t-h}^{-}$ are shown in (8).
@@
[…]
@@ 4.1 The Path-dependency property of realized volatility
In order to argue that the Chinese stock market is also characterized by path-dependence, we compare a path-dependence model that includes the historical price paths of stocks with a non-path-dependence model that assumes that price movements are independent of past paths. The data used are from the Wind database. Considering the non-negligible microstructure noise in the commodity market (Aït-Sahalia et al., 2005; Zhang et al., 2005), five-minute high-frequency data are selected as suggested by Corsi et al. (2008). The datasets include the SSE (000001), CSI 300 (000300), SSE 50 (000016), GEI (399006), SSE 50 ETF (510050), and STAR 50 (000688). The SSE covers the period from January 4, 2005, to December 29, 2023. The SSE 50 covers the same period. The CSI 300 covers the period from April 8, 2005, to December 29, 2023. The GEI covers the period from June 1, 2010, to December 29, 2023. The SSE 50 ETF covers the period from February 23, 2005, to December 29, 2023. The STAR 50 covers the period from July 23, 2020, to December 29, 2023. The realized volatility model based on the path dependence property is constructed as follows:
@@ 4.2 Contribution of Path-Dependence to Volatility
[…]
@@
[…]
@@
RV $0.201$ $0.399$ $7.857$ $95.914$ $7.175$ $0.007$ $-8.798^{***}$ $7873.7^{***}$
CJ $0.849$ $0.043$ $16.654$ $369.009$ $1.131$ $0.000$ $-12.424^{***}$ $137.6^{***}$
CV $0.142$ $0.243$ $5.828$ $55.344$ $3.898$ $0.005$ $-7.314^{***}$ $11988^{***}$
$\mathrm{RS}^{+}$ $0.099$ $0.223$ $14.884$ $345.431$ $6.543$ $0.002$ $-10.206^{***}$ $399.7^{***}$
$\mathrm{RS}^{-}$ $0.106$ $0.248$ $11.156$ $220.296$ $7.079$ $0.000$ $-8.979^{***}$ $3803.4^{***}$
$\mathrm{REX}^{+}$ $0.035$ $0.076$ $6.872$ $76.892$ $1.447$ $0.000$ $-7.619^{***}$ $6809.6^{***}$
$\mathrm{REX}^{-}$ $0.038$ $0.083$ $8.627$ $111.056$ $1.546$ $0.000$ $-9.561^{***}$ $6320.3^{***}$
$\mathrm{REX}^{m}$ $0.077$ $0.132$ $6.379$ $71.830$ $2.556$ $0.000$ $-7.341^{***}$ $12713^{***}$
$\mathrm{REXQ}^{+}$ $0.041$ $0.081$ $8.549$ $114.839$ $0.002$ $0.000$ $-8.857^{***}$ $5057.1^{***}$
$\mathrm{REXQ}^{-}$ $0.039$ $0.079$ $6.930$ $72.924$ $1.45$ $0.000$ $-7.487^{***}$ $4623.4^{***}$
$\mathrm{REXQ}^{m}$ $0.069$ $0.117$ $6.054$ $60.071$ $2.123$ $0.000$ $-7.194^{***}$ $11575^{***}$
@@
    Note: CV and CJ denote the continuous and jump components, respectively, derived from the underlying volatility model. The Augmented Dickey-Fuller (ADF) test examines the presence of a unit root in the series, with the null hypothesis positing that the series may be non-stationary. The Ljung-Box(10) test evaluates whether time series data exhibit autocorrelation. For ease of presentation, the mean, standard deviation, and maximum values are displayed after multiplying the original values by $10^{3}$. The symbols *, **∗ ∗, and ***∗ ∗ ∗ indicate rejection of the null hypothesis at the $10\%$, $5\%$, and $1\%$ significance levels, respectively.

### v2

@@ Abstract
Financial market volatility prediction has been a research focus in recent years. This paper presents a new volatility prediction framework that captures volatility and trend features by leveraging long - and short - term memory in price data. It integrates the heterogeneous autoregressive (HAR) model family to develop a path-dependent heterogeneous autoregressive model family (HAR-PD). And, we propose a HAR-REQ model based on the empirical quantile as a threshold, which exhibits stronger forecasting ability compared to the HAR-REX model. Subsequently, the predictive performance of the HAR-PD model family is evaluated by statistical tests using data from the Chinese stock market and compared with the benchmark HAR model family. The empirical results show that the HAR-PD model family has higher forecasting accuracy compared to the underlying HAR model family. In addition, robustness tests confirm the significant predictive power of the HAR-PD model family.
@@ 1 Introduction
By leveraging the entire historical trajectory of the underlying asset prices, the PDV model can more comprehensively capture long-term memory effects and complex dynamics, and remain highly sensitive to new market environments. The PDV model maintains forecast accuracy at different time scales and has a high sensitivity to extreme market events, which is particularly important for predicting potential large fluctuations. Parent (2022) integrated the concept of path dependence with the rough Heston model (Euch and Rosenbaum, 2018) to introduce the rough path-dependent volatility model. This model can capture a stronger Zumbach effect 11endnote: 1The strong Zumbach effect refers to the phenomenon where the distribution of future realized volatility depends not only on past squared returns but also on the entire trajectory of past realized volatility. The weak Zumbach effect refers to the covariance between past squared returns and future realized volatility (within a given duration) being greater than the covariance between past realized volatility and future squared returns.. These research findings not only enrich the theoretical foundation of PDV models but also open up new avenues for their practical application in finance.
@@
    Aiming to enhance predictive accuracy, we have refined the HAR-REX model by substituting the conventional cumulative normal distribution threshold with an innovative approach that employs empirical quantiles to re-segment realized volatility. Our empirical analyses reveal that the revamped HAR-REQ model, along with its extension, the HAR-PD-REQ model, exhibit significantly improved forecasting efficacy across daily, weekly, and monthly volatility horizons.
@@
The remainder of this paper is structured as follows. In Section 2, we introduce the fundamental theories of volatility and the construction of volatility forecasting models. In Section 3, we present all the benchmark models used in this study and the methods for combining models. In Section 4, we discuss the volatility of the Chinese financial market from a path-dependent perspective and conduct an empirical analysis of the dataset. We use various statistical testing methods, including the Model Confidence Set (MCS) test (Hansen et al., 2011) and the out-of-sample $R^{2}$ test (Campbell and Thompson, 2008), to evaluate the predictive capabilities of the HAR model family and the HAR-PD model family, and we report and discuss the results. In Section 5, we present the conclusion that the proposed HAR-PD model family demonstrates superior predictive ability.
@@ 2 Theoretical framework
$\mathrm{RV}_{t}=\sum_{i=1}^{n}r_{t,i}^{2}\xrightarrow{p}\int_{0}^{t}\sigma_{s}^{2}\,\mathrm{d}s+\sum_{0<s\leqslant t}\kappa_{s}^{2}\,$, (2)
@@ 2.1 The PDV model
The PDV model assumes that the asset price $S_{t}$ follows a functional SDE 22endnote: 2The well-posedness of (3) has been extensively studied and is beyond the scope of this paper; interested readers may refer to Ning and Wu (2021) and Cozma and Reisinger (2018):
@@
where $\mu_{t}$ denotes a progressively measurable stochastic process, $W_{t}$ is a standard Brownian motion, and $S_{u}(u\leq t)$ is the trajectory of $S_{u}$ for 0 to $t$, representing a functional mapping from the sample space $\mathbb{R}$ to the path space $\mathcal{C}(\mathbb{R})$. In contrast to classical jump-diffusion model (1), the PDV models (3) postulate that the volatility is dependent upon the entire trajectory of asset $S_{u}(u\leq t)$ instead of $S_{t}$. This implies that the volatility is determined not only by the current price level but also by the path through which the price reaches its current level. The path-dependent variable $G_{t}$ can incorporate various metrics, including the moving average ($S^{\Delta}_{t}$), exponentially weighted moments ($Y^{\Delta}_{t}$), and historical maximum or minimum prices ($Max^{\Delta}_{t}$ and $Min^{\Delta}_{t}$), where:
$S^{\Delta}_{t}$ $=\frac{\int_{0}^{\Delta}w_{\tau}S_{t-\tau}d\tau}{\int_{0}^{\Delta}w_{\tau}d\tau}\,,\quad Y^{\Delta}_{t}=\int_{-\infty}^{\Delta}\lambda e^{-\lambda(t-s)}\log\big\{S_{t}/S_{s}\cdot e^{-r(t-s)}\big\}ds,$
@@
where $K_{\lambda_{1}}$ is a kernel function that assigns different weights to past daily returns based on the lag term tit-i. Moreover, $K_{\lambda_{1}}$ is specified as exponential kernels:
@@ 2.2.1 Continuous and discrete components
where $\mu_{1}=\sqrt{2/\pi}$, and the factor n/(n1)n/(n-1) is used to adjust for sample size. Barndorff-Nielsen and Shephard (2003) demonstrated that under the assumption that the price process $S_{t}$ is a semimartingale, $\mathrm{RBV}_{t}$ serves as a consistent estimator of the integrated volatility and the difference between ${\rm RV}_{t}$ and $\mathrm{RBV}_{t}$ provides an estimator for jump variation.
$\mathrm{RBV}_{t}\xrightarrow{p}\int_{0}^{t}\sigma_{s}^{2}\,\mathrm{d}s\,,~~~{\rm and}~~~\mathrm{RV}_{t}-\mathrm{RBV}_{t}\xrightarrow{p}\sum_{0<s\leqslant t}\kappa_{s}^{2}\,.$
@@ 3.1 Benchmark: HAR model family
The model comprises three explanatory variables: $\mathrm{{RV}}_{t-1}$, $\mathrm{\overline{RV}}_{t-5}$, and $\mathrm{\overline{RV}}_{t-22}$, which correspond to short-term (daily), medium-term (weekly), and long-term (monthly) volatility, respectively. The coefficients $\beta_{0},\beta_{1}$, $\beta_{2}$, $\beta_{3}$ in (7) can be estimated using the loss function MSE for estimation. The lagged terms $\mathrm{\overline{RV}}_{t-5}$ and $\mathrm{\overline{RV}}_{t-22}$ represent five-step and twenty-two-step lags, respectively, and can be generalized as $\mathrm{\overline{RV}}_{t-h}$, where $h$ denotes the number of lag terms, calculated as:
@@
[…]
@@
[…]
@@ 3.2 Encoding path-dependent: HAR-PD models
$\mathrm{RV}_{t}=\beta_{0}+\beta_{1}{R}_{2,t-1}+\beta_{2}{\overline{R}}_{2,t-5}+\beta_{3}{\overline{R}}_{2,t-22}+\epsilon_{t}\,$, (11)
where ${R}_{2,t-1}$, ${\overline{R}}_{2,t-5}$, and ${\overline{R}}_{2,t-22}$ represent volatility features based on the moments t1{t-1}, t5{t-5}, and t22{t-22}, respectively. The construction method for ${\overline{R}}_{2,t-h}$ is similar (5).
@@
In the HAR-PD-CJ model, the volatility characteristics incorporate both jump and continuous components. The volatility characteristics describe the overall fluctuation within a specific time period, while jumps in financial markets indicate significant changes in asset prices. By incorporating jump characteristics, volatility predictions can be improved, particularly for fluctuations caused by public events, announcements, and other market uncertainties. Here, $\mathrm{{PDCJ}}_{t-1}$, $\mathrm{\overline{{PDCJ}}}_{t-5}$, and $\mathrm{\overline{{PDCJ}}}_{t-22}$ represent the jump components based on the path dependence at time t1t-1, t5t-5, t22t-22 respectively. Combined with the constructor of (4), the $\mathrm{PDCJ}_{t}$ can be constructed as:
@@
The benchmark HAR-RS model decomposes the realized semivariance into positive and negative components for volatility prediction. This decomposition captures both the actual price skewness within specific time periods and the dynamic evolution of realized skewness, thereby providing insights into volatility patterns. The construction method for $\mathrm{PDRS}_{t}$ is shown in (13), and the construction methods for $\mathrm{\overline{{PDRS}}}_{t-h}^{+}$ and $\mathrm{\overline{{PDRS}}}_{t-h}^{-}$ are shown in (8).
@@
The construction methods of $\mathrm{{PDREQ}}_{t-1}^{+}$, $\mathrm{\overline{{PDREQ}}}_{t-5}^{+}$, and $\mathrm{\overline{{PDREQ}}}_{t-22}^{+}$ are similar to those of (8) and (13).
@@ 4.1 The Path-dependency property of realized volatility
In order to argue that the Chinese stock market is also characterized by path-dependence, we compare a path-dependence model that includes the historical price paths of stocks with a non-path-dependence model that assumes that price movements are independent of past paths. The data used are from the Wind database. Considering the non-negligible microstructure noise in the commodity market (Aït-Sahalia et al., 2005; Zhang et al., 2005), five-minute high-frequency data are selected as suggested by Corsi et al. (2008). The datasets 33endnote: 3The data that support the findings of this study are available from the corresponding author upon reasonable request. include the SSE (000001), CSI 300 (000300), SSE 50 (000016), GEI (399006), SSE 50 ETF (510050), and STAR 50 (000688). The SSE covers the period from January 4, 2005, to December 29, 2023. The SSE 50 covers the same period. The CSI 300 covers the period from April 8, 2005, to December 29, 2023. The GEI covers the period from June 1, 2010, to December 29, 2023. The SSE 50 ETF covers the period from February 23, 2005, to December 29, 2023. The STAR 50 covers the period from July 23, 2020, to December 29, 2023. The realized volatility model based on the path dependence property is constructed as follows:
@@ 4.2 Contribution of Path-Dependence to Volatility
(a) SSE residuals ACF under models $M_{1}$$M_{5}$
(b) CSI 300 residuals ACF under models $M_{1}$$M_{5}$
(c) SSE 50 residuals ACF under models $M_{1}$$M_{5}$
(d) GEI residuals ACF under models $M_{1}$$M_{5}$
(e) SSE 50 ETF residuals ACF under models $M_{1}$$M_{5}$
(f) STAR 50 residuals ACF under models $M_{1}$$M_{5}$
Figure 5: The residuals ACF of each index under model $M_{1}$$M_{5}$
@@
As shown in Figure 5, it shows the results of the residual autocorrelation of each sample data in the model $M_{1}$$M_{5}$ with a lag of 500 items. And the models $M_{1}$$M_{5}$ are represented by green, red, blue, magenta, and cyan respectively and the two horizontal lines represent confidence intervals. Taking Figure 5(a) as an example, Model $M_{3}$ exhibits strong and persistent autocorrelation in residuals, indicating the presence of unexplained dynamic features in the data, such as nonlinear characteristics, or other factors not captured by the model. In contrast, Model $M_{2}$ significantly reduces residual autocorrelation, indicating that the model effectively captures the primary impact of lagged variables on the target variable. As shown in Figures 5(b)5(e), model $M_{2}$ exhibits a low degree of autocorrelation, with most values falling within the confidence interval. This result indicates that the model $M_{2}$ has effectively reduced the unexplained information in the residuals, enhancing overall fitting accuracy and explanatory power. This analysis further confirms that in modeling path-dependent, volatility features play a dominant role in explaining volatility dynamics. This conclusion is consistent with the findings of Parent (2023). Table 2 shows the fitting performance of Models $M_{1}$$M_{5}$ on different datasets. The results indicate that Model $M_{2}$ has significantly superior fitting performance on all samples (both train and test) compared with other models. By contrast, Model $M_{3}$ has relatively poor fitting performance. The above illustrates the contribution of the volatility feature to explaining realized volatility, and we use this feature to construct the HAR-PD-RV model at (11) and use the structure of Model $M_{2}$ to construct the other HAR-PD models.
@@
RV $0.201$ $0.399$ $7.857$ $95.914$ $7.175$ $0.007$ $8.798^{***}$ $7873.7^{***}$
CJ $0.849$ $0.043$ $16.654$ $369.009$ $1.131$ $0.000$ $12.424^{***}$ $137.6^{***}$
CV $0.142$ $0.243$ $5.828$ $55.344$ $3.898$ $0.005$ $7.314^{***}$ $11988^{***}$
$\mathrm{RS}^{+}$ $0.099$ $0.223$ $14.884$ $345.431$ $6.543$ $0.002$ $10.206^{***}$ $399.7^{***}$
$\mathrm{RS}^{-}$ $0.106$ $0.248$ $11.156$ $220.296$ $7.079$ $0.000$ $8.979^{***}$ $3803.4^{***}$
$\mathrm{REX}^{+}$ $0.035$ $0.076$ $6.872$ $76.892$ $1.447$ $0.000$ $7.619^{***}$ $6809.6^{***}$
$\mathrm{REX}^{-}$ $0.038$ $0.083$ $8.627$ $111.056$ $1.546$ $0.000$ $9.561^{***}$ $6320.3^{***}$
$\mathrm{REX}^{m}$ $0.077$ $0.132$ $6.379$ $71.830$ $2.556$ $0.000$ $7.341^{***}$ $12713^{***}$
$\mathrm{REXQ}^{+}$ $0.041$ $0.081$ $8.549$ $114.839$ $0.002$ $0.000$ $8.857^{***}$ $5057.1^{***}$
$\mathrm{REXQ}^{-}$ $0.039$ $0.079$ $6.930$ $72.924$ $1.45$ $0.000$ $7.487^{***}$ $4623.4^{***}$
$\mathrm{REXQ}^{m}$ $0.069$ $0.117$ $6.054$ $60.071$ $2.123$ $0.000$ $7.194^{***}$ $11575^{***}$
@@
    Note: CV and CJ denote the continuous and jump components, respectively, derived from the underlying volatility model. The Augmented Dickey-Fuller (ADF) test examines the presence of a unit root in the series, with the null hypothesis positing that the series may be non-stationary. The Ljung-Box(10) test evaluates whether time series data exhibit autocorrelation. For ease of presentation, the mean, standard deviation, and maximum values are displayed after multiplying the original values by $10^{3}$. The symbols *, **, and *** indicate rejection of the null hypothesis at the $10\%$, $5\%$, and $1\%$ significance levels, respectively.